OpenStudy (helpmeplease!!!):

the tongue of a chameleon is one and a half times its body length write this as a decimal

OpenStudy (ddcamp):

\[1 + \frac{ 1 }{ 2 } = 1 + 0.5 = 1.5\]

OpenStudy (helpmeplease!!!):

Thank you.